# Qualification-free convex analysis via the joint supporting subspace

### v3

@@ 3 Preliminaries
For any set $S\subseteq C$, we denote by $F_{C}(S)$ the face of $C$ generated by $S$; it is the smallest face of $C$ that contains $S$ (it is also the intersection of all faces containing $S$, which, it can be shown, is itself a face).
@@
if $x\in C$, and $N_{C}(x)=\emptyset$ otherwise.
We also recall some objects from convex analysis. Let $\bar{\mathbb{R}}=\mathbb{R}\cup\{-\infty,\infty\}$. For a convex function $f:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, the subdifferential is
@@ 4 The joint supporting subspace
Now that the joint supporting subspace has been thoroughly identified, we move on to describing its core property. Recall that for two convex functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, a common qualification condition is that for convex sets $C:=\operatorname{dom}(f),D:=\operatorname{dom}(g)$,
@@ 5 Main results
The next result implies that when standard qualification conditions hold, the classical formulation of our generalizations is trivially recovered.
@@ 5.1 Subdifferential calculus
Let $g\in\mathbb{R}^{m}\to\bar{\mathbb{R}}$ be a proper convex function, and $A\in\mathbb{R}^{m\times n}$ a matrix. Then for $T:=T(\operatorname{dom}(g),\operatorname{\mathrm{range}}(A))$,
@@ 5.3 Convex optimization
which is a weakening of Slater’s condition. A program with this property is called strongly consistent [rockafellarConvexAnalysis1970, p. 300].
@@
By Corollary 4.1, $(A\operatorname{dom}(f))\cap\operatorname{dom}(g)\subseteq T_{a}$, and therefore $\iota_{T_{a}}(Ax)=0$ for all $x\in\operatorname{dom}(f)\cap A^{-1}\operatorname{dom}(g)$. Additionally,
@@
$=$ $\operatorname{relint}((A\operatorname{dom}(f))\cap T_{a})\cap\operatorname{relint}(\operatorname{dom}(g)\cap T_{a})\neq\emptyset.$ (10)
where Equation 8 holds by [rockafellarConvexAnalysis1970, Theorem 6.6]; Equation 10 uses the identity that $A(S_{1}\cap A^{-1}(S_{2}))=AS_{1}\cap S_{2}$ for any sets $S_{1}\subseteq\mathbb{R}^{n}$, $S_{2}\subseteq\mathbb{R}^{m}$; and the last inequality is guaranteed by Corollary 4.5. ∎
@@ 6 A framework for local face lattices
1), $(\implies)$ we show that if $L$ is a supporting subspace, then $F:=C\cap L$ satisfies both conditions of Theorem 6.1:
@@
2), ($\implies$) If $F$ is a face, then we argue that $\operatorname{aff}(F)$ is a suitable supporting subspace. Indeed, by Theorem 6.1 we know that $F=C\cap\operatorname{aff}(F)$ and that $C\setminus F$ is convex. Note that $C\setminus F=C\setminus(C\cap\operatorname{aff}(F))=C\setminus\operatorname{aff}(F)$, and therefore $C\setminus\operatorname{aff}(F)$ is convex, while $F=C\setminus\operatorname{aff}(F)$, which shows that $\operatorname{aff}(F)$ is a supporting subspace that selects $F$ by intersection.
@@
We can see that a normal cone is a dual object which “describes” the original cone only up to its closure.
We require a dual object which has sufficient descriptive power to go beyond the closure, by encoding information about the boundary. To this end, we introduce the notion of a nested normal cone. Let $C\subseteq\mathbb{R}^{n}$ be convex and $U\subseteq\mathbb{R}^{n}$ be a subspace, then $\forall x\in\mathbb{R}^{n}$,
$N_{C}^{U}(x):=N_{C\cap U}(x)\cap U.$
@@
The hyperplanes corresponding to the vectors in this normal cone are supporting hyperplanes of $C$ that contain $S$. Note also that $N_{C}(S)=\bigcap_{s\in S}N_{C}(s).$ The idea of normal cones that are located at sets appears in the earlier work [lu_NormalFansPolyhedral_2008], where normal cones are at faces of polytopes, instead at points. The normal cone at a set has an elegant connection to the notion of relative interior of that set: the normal cone at any point in the relative interior of a convex set $S\subseteq C$ will match the normal cone of $S$.
Combining the two generalizations, we get the following.
@@
is a nested normal cone of $C$ at $S$, and we label any $v\in N_{C}^{U}(S)$ to be a nested normal of $C$ at $S$ in $U$.
The idea of normal cone defined at a set has an implicit relation with the notion of relative interior.
@@
Let $C\subseteq\mathbb{R}^{n}$ be convex and $S\subseteq C$ a non-empty convex set. Let $U\subseteq\mathbb{R}^{n}$ be a subspace, and $x\in\operatorname{relint}(S)$. Then
@@
Let $y\in N_{C}^{U}(S)$. By Definition 6.3, $\forall z,x\in S,$ since $x\in S$ and $z\in C$, $\langle x-z,y\rangle\leq 0,$ and since $x\in C$ and $z\in S$ $\langle x-z,y\rangle\geq 0$. Then $y\perp S-S$, which shows the inclusion.
@@ 7.1 Regularity in the joint supporting subspace
Thus, for all $y\in C$ and $s\in S$, $\langle v,y-s\rangle\geq 0$, so $v\in N_{C}(S)$. Similarly, $v\in N_{D}(S)$, whence $v\in N_{C}(S)\cap(-N_{D}(S))\neq\{0\}$.\}\$. ∎

### v4

@@ Abstract
Keywords: facial reduction, subdifferential calculus, exact duality, qualification conditions, constraint qualifications, convex analysis
@@ 3 Preliminaries
For any set $S\subseteq C$, we denote by $F_{C}(S)$ the face of $C$ generated by $S$; it is the smallest face of $C$ that contains $S$. (It is also the intersection of all faces containing $S$, which, it can be shown, is itself a face.)
@@
if $x\in C$, and $N_{C}(x)=\emptyset$ otherwise. We also recall some objects from convex analysis. Let $\bar{\mathbb{R}}=\mathbb{R}\cup\{-\infty,\infty\}$. For a convex function $f:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, the subdifferential is
@@ 4 The joint supporting subspace
We defer the proof to Section 7.3. Now that the joint supporting subspace has been thoroughly identified, we move on to describing its core property. Recall that for two convex functions $f,g:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, a common qualification condition is that for convex sets $C:=\operatorname{dom}(f),D:=\operatorname{dom}(g)$,
@@ 5 Main results
The next result implies that when standard qualification conditions hold, the classical formulations of our generalizations are trivially recovered.
@@ 5.1 Subdifferential calculus
Let $g:\mathbb{R}^{m}\to\bar{\mathbb{R}}$ be a proper convex function, and $A\in\mathbb{R}^{m\times n}$ a matrix. Then for $T:=T(\operatorname{dom}(g),\operatorname{\mathrm{range}}(A))$,
@@ 5.3 Convex optimization
which is a weakening of Slater’s condition. A program with this property is called strongly consistent [rockafellarConvexAnalysis1970, Theorem 31.2].
@@
By Corollary 4.1, $A\operatorname{dom}(f)\cap\operatorname{dom}(g)\subseteq T_{a}$, and therefore $\iota_{T_{a}}(Ax)=0$ for all $x\in\operatorname{dom}(f)\cap A^{-1}\operatorname{dom}(g)$. Additionally,
@@
$=$ $\operatorname{relint}(A\operatorname{dom}(f)\cap T_{a})\cap\operatorname{relint}(\operatorname{dom}(g)\cap T_{a})\neq\emptyset.$ (10)
where Equation 8 holds by [rockafellarConvexAnalysis1970, Theorem 6.6]; Equation 10 uses the identity that $A(S_{1}\cap A^{-1}(S_{2}))=AS_{1}\cap S_{2}$ for any sets $S_{1}\subseteq\mathbb{R}^{n}$, $S_{2}\subseteq\mathbb{R}^{m}$; and that the set in Equation 10 is non-empty is guaranteed by Corollary 4.5. ∎
@@ 6 A framework for local face lattices
1) ($\implies$) We show that if $L$ is a supporting subspace, then $F:=C\cap L$ satisfies both conditions of Theorem 6.1:
@@
2) ($\implies$) If $F$ is a face, then we argue that $\operatorname{aff}(F)$ is a suitable supporting subspace. Indeed, by Theorem 6.1 we know that $F=C\cap\operatorname{aff}(F)$ and that $C\setminus F$ is convex. Note that $C\setminus F=C\setminus(C\cap\operatorname{aff}(F))=C\setminus\operatorname{aff}(F)$, and therefore $C\setminus\operatorname{aff}(F)$ is convex, while $F=C\setminus\operatorname{aff}(F)$, which shows that $\operatorname{aff}(F)$ is a supporting subspace that selects $F$ by intersection.
@@
We can see that a normal cone is a dual object which “describes” the original cone only up to its closure. We require a dual object which has sufficient descriptive power to go beyond the closure, by encoding information about the boundary. To this end, we introduce the notion of a nested normal cone. Let $C\subseteq\mathbb{R}^{n}$ be convex and $U\subseteq\mathbb{R}^{n}$ be a subspace, then $\forall x\in\mathbb{R}^{n}$,
$N_{C}^{U}(x):=N_{C\cap(x+U)}(x)\cap U.$
@@
The hyperplanes corresponding to the vectors in this normal cone are supporting hyperplanes of $C$ that contain $S$. Note also that $N_{C}(S)=\bigcap_{s\in S}N_{C}(s).$ The idea of normal cones that are located at sets appears in the earlier work [lu_NormalFansPolyhedral_2008], where normal cones are defined for faces of polytopes, instead of at points. Combining the two generalizations, we get the following.
@@
is a nested normal cone of $C$ at $S$, and we label any $v\in N_{C}^{U}(S)$ a nested normal of $C$ at $S$ in $U$.
The normal cone at a set has an elegant connection to the notion of relative interior of that set: the normal cone at points in the relative interior match the normal cone of the set.
@@
Let $C\subseteq\mathbb{R}^{n}$ be convex and $S\subseteq C$ a non-empty convex set. Let $U\subseteq\mathbb{R}^{n}$ be a subspace such that $S-S\subseteq U$, and $x\in\operatorname{relint}(S)$. Then
@@
Let $y\in N_{C}^{U}(S)$. By Definition 6.3, $\forall z,x\in S,$ since $x\in C$ and $z\in S$, $\langle x-z,y\rangle\leq 0,$ and since $x\in S$ and $z\in C$, $\langle x-z,y\rangle\geq 0$. Then $y\perp S-S$, which shows the inclusion.
@@ 7.1 Regularity in the joint supporting subspace
Thus, for all $y\in C$ and $s\in S$, $\langle v,y-s\rangle\geq 0$, so $v\in N_{C}(S)$. Similarly, $v\in N_{D}(S)$, whence $v\in N_{C}(S)\cap(-N_{D}(S))\neq\{0\}$. ∎
@@ 7.3 Second characterization
Proof of Corollary 4.4.
We consider the formula
$T_{i+1}=T_{i}\cap(N^{T_{i}}_{C}(x)\cap[-N^{T_{i}}_{D}(x)])^{\perp}.$
from Theorem 7.1. We show that the r.h.s. is the same set as
$T_{i}\cap(N_{C\cap(T_{i}+x)}(x)\cap[-N_{D\cap(T_{i}+x)}(x)])^{\perp}.$
We use a substitution as in Lemma 7.2 with $V=\mathbb{R}^{n}$.
$T_{i}\cap(N_{C\cap(T_{i}+x)}(x)\cap[-N_{D\cap(T_{i}+x)}(x)])^{\perp}$
$=$ $T_{i}\cap((N^{T_{i}}_{C}(x)+T_{i}^{\perp})\cap[-N^{T_{i}}_{D}(x)+T_{i}^{\perp}])^{\perp}$
$=$ $T_{i}\cap(N^{T_{i}}_{C}(x)\cap[-N^{T_{i}}_{D}(x)]+T_{i}^{\perp})^{\perp}$
$=$ $T_{i}\cap(N^{T_{i}}_{C}(x)\cap[-N^{T_{i}}_{D}(x)])^{\perp}.$
The last equality holds because $N^{T_{i}}_{C}(x)\cap[-N^{T_{i}}_{D}(x)]\subseteq T_{i}$, so we may use the identity that $T_{i}\cap(S+T_{i}^{\perp})^{\perp}=T_{i}\cap S^{\perp}$ for any set $S\subseteq T_{i}$. ∎
@@ 7.4 Acknowledgements
Declarations
The author is supported by the University of British Columbia. The author has no relevant financial or non-financial interests to disclose.